OpenStudy (linn99123):

A square pool has an area of 292 square feet. To the nearest tenth, what is the length of one side of the pool? A. 8.5 feet B. 17.1 feet C. 73.0 feet D. 146.0 feet

OpenStudy (bibby):

Work backwards. What is the area of a square?

OpenStudy (linn99123):

292

OpenStudy (bibby):

A square. Any square. This square

OpenStudy (linn99123):

im confused

OpenStudy (bibby):

k I'll draw the problem out then. If you recall, the area of a square is x*x where x is one of the sides

OpenStudy (bibby):

So we have: Area=x*x=x^2 area=292 x^2=292 Can you solve

OpenStudy (linn99123):

so i divide it into 4?

OpenStudy (linn99123):

its 73

OpenStudy (bibby):

\[x^{2}=292\] The opposite operation of squaring something is taking the root of that. So you have to take the square root of 292

OpenStudy (anonymous):

was he answer